# ECLeKTic: a Novel Challenge Set for Evaluation of Cross-Lingual Knowledge Transfer

### v2

@@ 4.2 Shared Script Eases Transfer
In order to provide further insight into the factors that affect knowledge transfer, Figure 3 details a breakdown per language of our results in terms of transfer score of our best performing model, Gemini 2.0 Pro. This analysis shows that the model’s ability to transfer knowledge is highly dependent on the source and target language, with average scores ranging from 23.5 (transfer from Portuguese to Japanese) to 100.0 (from German to Indonesian).
It is also possible to see that transfer is much higher between languages with the same script. This is evident from the high transfer score between German, English, French, Spanish, Italian, Portuguese, and Indonesian. Note that the latter is not genealogically related to the rest of the Latin-written languages, yet it performs on par with the others when serving as a source or a target to other Latin-written languages. The dependence on script can also be seen in the transfer scores between Chinese and Japanese, especially when compared to the transfer from Japanese to other languages. This finding aligns with previous works in the literature on the importance of script to transfer Malkin et al. (2022); Mittal et al. (2023); Ifergan et al. (2024).
Additionally, this analysis reveals an asymmetry in transfer scores depending on the role of the language, as a source or a target. For example, knowledge seems to be easily transferable from Hindi, mostly to Latin-written languages and to Chinese, resulting in a macro-averaged transfer score of 78.6. But the transfer to Hindi is much worse, averaging in only 59.6 when Hindi is the target language.
A similar breakdown for the overall success metric can be found in Appendix B.
4.3 Bigger Isn’t Necessarily Better
@@
The difference in the curves of the transfer and the overall scores is clearly evident. The performance in terms of overall score continues to improve with almost every increase in model size, and it seems plausible that an even bigger version of Qwen would have an even better performance. On the other hand, in terms of transfer score, the performance improves rapidly with the increase in model size up until 14B parameters and then somewhat saturates, as more than a 5-fold increase in model size only give an improvement of about 7 percentage points in transfer. Taken together, this leads to the conclusion that the improvement in the overall scores of the bigger model comes mostly from their ability to retrieve more facts in their source language while the proportion of facts transferred rises only marginally.
@@ 4.4 Hinting LLMs Into Success Is Not Easy
In the evaluation setting we examined until now — a zero-shot setting where models are only given the question itself with no instruction — we found that models have a significant room for improvement. In order to characterize the pitfalls that make models fail and how to avoid them, we conducted an ablation study by incorporating more and more information into the prompt given to the model. We experimented with Gemini 2.0 Flash in the following settings:
@@
Source language name. This setting is very similar to the previous one but the name of language with the relevant knowledge is given explicitly as part of the prompt.
@@
The prompts used in all of these settings are given in Appendix A.
The results, given in Table 3, show that while simply hinting the model towards cross-lingual knowledge is not enough and provides only insignificant improvement. Revealing the source language and the topic leads to a much improved performance of about 7 and 12 percentage points, respectively, indicating that in some cases the knowledge may be partially available to the model, but it requires some guidance.
@@
5 ECLeKTic Popular Pages
| Model | Popular Pages | |
| Overall | Transfer | |
| Gemini 2.0 Pro | 36.71.2 | 67.81.7 |
| GPT 4o | 34.51.2 | 67.31.7 |
| Gemini 2.0 Flash | 31.61.2 | 65.81.8 |
| Claude 3.5 Sonnet | 23.81.1 | 56.82.0 |
| Gemma 2 9B | 6.00.6 | 33.02.9 |
| Mistral Nemo | 4.00.5 | 31.53.4 |
| Qwen 2.5 7B | 1.50.3 | 18.93.6 |
| Olmo 2 7B | 0.80.2 | 9.32.5 |
Beyond reliance on Wikipedia article distribution, when constructing ECLeKTic we also made a decision to base the questions and answers on pages that exist in only one language out of the 12 in our selection. While providing control over the language from which the knowledge has to be transferred, it also means that ECLeKTic may include many questions on topics that are somewhat marginal and less consequential to users in the target languages.
To examine whether our results are due to limited exposure of the models to their topics, we experimented with another variant of our dataset, namely ECLeKTic popular pages. In this version we sampled articles that were popular in terms of views, instead of prioritizing lack of equivalents. Thus, this version gives emphasis to topics that are more likely to interest average users and appear more frequently outside Wikipedia.
Specifically, from each Wikipedia we took the 200 articles that had the most views during April 2023 and lacked equivalent in at least one of the 12 Wikipedias. Then,S through the same pipeline described in Section 3.1, we created, translated and verified questions and answers based on the content of the articles. We ended up with 964 unique question/answer pairs. However, each example was evaluated on a subset of target languages consisting only of languages whose Wikipedias do not include an article on the question’s topic, resulting in a total of 6,628 examples summed over all languages.111111The data for this variant is also included in https://www.kaggle.com/datasets/googleai/eclektic.
We evaluated the same models of Section 4.1 on this data as well. The results, detailed in Table 4, are inline with the results over the main ECLeKTic dataset in Table 2. The order of models by performance didn’t change, Gemini 2.0 Pro is the best model, followed by GPT 4o, and open 7-9B models far worse. This experiment indicates on the robustness of our results to the article selection criterion and on the validity of our assumptions about the link between Wikipedia articles distribution and the necessity of transfer.
6 Related Work and Discussions
@@ 6.1 Types of Cross-Lingual Transfer
The first distinction in the literature concerns what is being transferred, specifically the difference between cross-lingual skill transfer and knowledge transfer (Rajaee and Monz, 2024). Cross-lingual skill transfer is the ability to generalize a given skill to unseen languages regardless of the language that was used to learn it. E.g. perfecting multilingual summarization when learning to summarize from English data, or excelling at multilingual instruction following while only exposed to a few languages (Hu et al., 2020; Turc et al., 2021; Malkin et al., 2022; Huang et al., 2023; Shaham et al., 2024). On the other hand, cross-lingual knowledge transfer is the ability to retrieve factual knowledge from one language’s data when queried with any language (Asai et al., 2020; Limkonchotiwat et al., 2022; Mittal et al., 2023; Chua et al., 2024; Litschko et al., 2024).
@@
In addition, methods for evaluation of cross-lingual transfer can also be orthogonally categorized into two broad approaches: fine-tuning on sources, then testing on targets (e.g., Shaham et al., 2024; Limkonchotiwat et al., 2022) vs. zero-shot evaluation of the targets (e.g., Malkin et al., 2022; Chua et al., 2024). While the earlier enables higher control over the data and experimental setting, the latter is less expensive and reflects how models behave in the wild.
@@
Lastly, when dealing with cross-lingual knowledge transfer, the knowledge can be either parametric (e.g., Rajaee and Monz, 2024) or contextual (e.g., Chua et al., 2024; Mondshine et al., 2025), a distinction explored by Neeman et al. (2022) in a monolingual setting.
@@
Within this taxonomy of cross-lingual transfer related works, ECLeKTic clearly belongs to methods evaluating parametric knowledge transfer in a zero-shot setting. In addition, the experiment in Section 4.4 gradually transition from parametric to contextual knowledge transfer, with the cross-lingual open-book experiment occupying the other end of that spectrum.
6.2 ECLeKTic and Cross-Lingual Consistency
Cross-lingual knowledge transfer is closely related to cross-lingual consistency, which has been explored extensively in recent literature. Starting with monolingual English settings, early works noticed that LLMs lack a guarantee on consistency due to the statistical nature of their training. That is to say that models may generate contradictory statements when presented with semantically equivalent inputs Kassner and Schütze (2020); Ravichander et al. (2020) or may rephrase identical factual information in different ways Elazar et al. (2021). In the context of modern LLMs, there has been increasing attention on ensuring output consistency when faced with variations in prompts Mizrahi et al. (2024); Sclar et al. (2024); Zhao et al. (2024b). A variety of methods have been proposed to improve this consistency, including the augmentation and diversification of instructions during training Liang et al. (2023); Zhao et al. (2024b).
In multilingual models, the challenge becomes even more complex, as it introduces the issue of cross-lingual inconsistency, where models fail to provide consistent responses to semantically equivalent inputs in different languages. To assess cross-lingual consistency, earlier works translated English datasets, either derived from knowledge bases Kassner et al. (2021); Jiang et al. (2020); Ifergan et al. (2024) or from NLP tasks Ohmer et al. (2023), into multiple languages. These studies then measured the variance in LLM responses, typically in terms of answer overlap or ranking Qi et al. (2023).
However, when applying translation across-the-board to entire datasets, these consistency-focused benchmarks may include some well-known facts that the model saw and memorize in many languages separately and therefore could easily predict consistently (see discussion in Section 2). Moreover, starting with English-constructed datasets may introduce biases into these benchmarks.
The QA task of ECLeKTic, although created with transfer in mind, may also serve as a better benchmark for cross-lingual consistency. To begin with, ECLeKTic is not generated in English but in all languages of the dataset. But more importantly, by targeting specific knowledge that is less known in most languages, we present models with a far greater challenge in keeping their answers consistent. It is therefore possible to view the results over ECLeKTic also as a tighter upper bound on the consistency of multilingual models.
@@ 7 Conclusions
In this paper we presented ECLeKTic, a closed-book QA dataset for evaluating the abilities of models to transfer knowledge from their parametric memory across languages. This black-box evaluation was made possible by carefully phrasing questions that target topics that are highly visible in one language and not in any of the others. Our benchmark allows for simple and reliable transfer evaluation which is, for the first time, easily applicable also to API-fenced models. Our results show that cross-lingual knowledge transfer is a difficult task that is far from being solved, and that ECLeKTic can be employed to indicate on the progress made towards consistent and inclusive multilingual language models.
@@ 8 Limitations
Although the data ECLeKTic covers varied languages, its coverage is obviously partial. It covers only 6 out of the 10 most spoken languages according to Ethnologue,121212https://www.ethnologue.com/statistics/ and 8 out of the 10 most active Wikipedias in terms of active users.131313https://en.wikipedia.org/wiki/List_of_Wikipedias However, due to the translation of all examples to all languages, adding significantly more languages may make the evaluation using ECLeKTic less cheap and fast.
@@
The limited number of languages also make the conclusions of Section 4.2 less unequivocal, but the fact the previous works pointed to the same conclusions Malkin et al. (2022); Mittal et al. (2023); Ifergan et al. (2024) provides some reassurance.
Appendix A Prompts
A.1 ECLeKTic Creation
@@
During the creation of ECLeKTic, LLMs were used for question and answer generation and for the translation of the examples from their respective source languages to all other target languages. Although human annotators verified and fixed their outputs, we also provide here the prompts that were used.
Question and Answer Generation
**Task:** Formulate a question in {lang_name} that requires a deep understanding of a given {lang_name} Wikipedia paragraph.
**Requirements:**
* **Context-Specific:** The question must be answerable solely through information presented within the paragraph, excluding general knowledge or common sense.
* **Self-Contained:** The question should be completely self-explanatory, providing all necessary context within its phrasing. Assume the reader has no access to the paragraph when answering the question.
* **Single Concrete Factual Detail:**
- The question should not require multiple answers or involve listing multiple details.
- Avoid asking about opinions, interpretations.
- In you can’t answer the question, prefer to generate another question.
- Focus on extracting a specific, concrete, factual detail that the paragraph directly states.
- Be specific:
- If you are asking about an entity be clear about it – Use full names for example.
- Mention expected granularity: If you are asking about a date, instead of asking "when", ask for a decade, year, month, date etc. If you are asking about a location, instead of asking "where", ask for a country, state, city, street, landmark etc.
- Avoid asking questions that their answers are acronyms.
Even for non-English examples keep the convention of using the English words "Paragraph", "Response", "question" and "answer" for specifying the parts being generated.
Generate only the question and answer. No need to continue with additional examples.
**Examples:**
Paragraph: The Great Barrier Reef is the world’s largest coral reef system, composed of over 2,900 individual reefs and 900 islands stretching for over 2,300 kilometers (1,400 mi) over an area of approximately 344,400 square kilometers (133,000 sq mi). The reef is located in the Coral Sea, off the coast of Queensland, Australia. The Great Barrier Reef can be seen from outer space and is the world’s biggest single structure made by living organisms.
Response: Question: Where is the Great Barrier Reef located? Answer: Coral Sea, off the coast of Queensland, Australia
Paragraph: Die Cazoo Snookerweltmeisterschaft 2023 wurde vom 15. April bis 1. Mai im Crucible Theatre in Sheffield ausgetragen. Mit ihr endete die Saison 2022/23 der World Snooker Tour.[1] Titelverteidiger Ronnie O’Sullivan scheiterte im Viertelfinale gegen Luca Brecel. Der Belgier erreichte das Finale und schlug dort den vierfachen Weltmeister Mark Selby mit 18:15. Brecel ist damit der erste Kontinentaleuropäer, der Weltmeister wurde. In diesem Jahr wurden noch weitere Bestmarken in Bezug auf die 47-jährige „Crucible-Ära“ aufgestellt. Unter anderem übertraf Ronnie O’Sullivan mit seiner 31. Endrundenteilnahme die 30 Teilnahmen von Steve Davis.[2] O’Sullivan erzielte auch sein 200. WM-Century-Break. Zweimal wurde ein Maximum Break erzielt, was es 2008 bereits einmal gegeben hatte; das „perfekte Break“ in einem WM-Finale gelang 2023 erstmals Mark Selby.
Response: Question: Gegen wen verlor Ronnie O’Sullivan im Viertelfinale der Snooker-Weltmeisterschaft 2023? Answer: Luca Brecel
Paragraph: {context}
Response:
Translation
@@
Translate the provided text from {in_lang} to {out_lang} while maintaining the original meaning and intent. Ensure accuracy and preserve the entities and concepts expressed in the source text.
Input in {in_lang}: {text} Response in {out_lang}:
A.2 Judging Prompt
**Task:** Determine if an answer to the question is supported by a given text.
**Input (in {target_language}):** - Text - Question - Answer
@@ A.2 Judging Prompt
**Single Word Output (in English):** - YES: Answer is derived from the text. - NO: Answer is not derived from the text.
Text: {context}
Question: {question}
Answer: {predicted_answer}
Output:
A.3 Prompts for Ablation in Section 4.4
Closed-book
{question}
General hint
Answer the following question based on your knowledge in another language.
{question}
Source language name
Answer the following question based on your knowledge in {in_lang}.
{question}
Source language title
@@
Answer the following question based on your knowledge in {in_lang} about {original_title}.
{question}
Cross-lingual open-book
Context: {original_context}
{question}
Appendix B Per-Language Breakdown - Overall Success

### v3

@@ 4.2 Shared Script Eases Transfer
In order to provide further insights into the factors affecting knowledge-transfer, Figure 5 details a per-language breakdown of our results in terms of transfer score of our best performing model, Gemini 2.0 Pro. This analysis shows that the model’s ability to transfer knowledge is highly dependent on the source and target language, with average scores ranging from 23.5% (transfer from Portuguese to Japanese) to 100% (from German to Indonesian).
Transfer is much higher between languages with the same script, as evident from high transfer score between German, English, French, Spanish, Italian, Portuguese and Indonesian. Note that the latter is not genealogically related to the rest of the Latin-written languages, yet it performs on par with the others when serving as a source or a target to other Latin-written languages. The dependence on script can be also seen in the transfer scores between Chinese and Japanese, especially when compared to the transfer of Japanese with other languages. This finding aligns with previous works in the literature on the importance of script to transfer (Malkin et al., 2022; Mittal et al., 2023; Ifergan et al., 2024).
Additionally, this analysis reveals an asymmetry in transfer scores depending on the role of the language, as a source or a target. For example, knowledge seems to be easily transferable from Hindi, mostly to Latin-written languages and Chinese, resulting in a macro-averaged transfer score of 78.6%. But the transfer to Hindi is much worse – only 59.6% when Hindi is the target language.
@@
A similar breakdown for the overall success metric can be found in Appendix C.
4.3 Bigger Isn’t Necessarily Better
@@
The difference in the curves of the transfer and the overall scores is evident. The performance in terms of overall score continues to improve with almost every increase in model size, and it seems plausible that an even bigger version of Qwen would have an even better performance. On the other hand, in terms of transfer score, the performance improves rapidly with the increase in model size up until 14B parameters and then somewhat saturates, as more than a 5-fold increase in model size only give an improvement of about 7 percentage points in transfer. Taken together, this leads to the conclusion that the improvement in the overall scores of the bigger model comes mostly from their ability to retrieve more facts in their source language while the proportion of facts transferred rises only marginally.
@@ 4.4 Hinting LLMs Into Success Is Not Easy
In the evaluation setting we examined until now – where models are only given the question itself – we found that models have a significant room for improvement. In order to characterize the pitfalls that make models fail and how to avoid them, we conducted an ablation study by incorporating more and more information into the prompt given to the model. We experimented with Gemini 2.0 Flash in the following settings:
@@
Source language name. This setting is very similar to the previous one but the name of the language with the relevant knowledge is given explicitly as part of the prompt.
@@
The prompts used in all of these settings are in Appendix B.
The results, in Table 2, show that simply hinting the model towards cross-lingual knowledge is not enough and provides only insignificant improvement. Revealing the source language and the topic leads to a much improved performance of about 7 and 12 percentage points, respectively, indicating that in some cases the knowledge is partially available to the model, but requires some guidance.
@@
5 Related Work
5.1 White-Boxing Methods for Transfer Evaluation
This paper presents a QA dataset to evaluate cross-lingual knowledge transfer in a black-box manner without access to the models’ inner workings. Another line of existing works aims to evaluate the same model behaviour by direct observation of the internal mechanisms of open models.
One line of works utilized knowledge editing to determine whether the model’s representation of knowledge is causally linked across languages (Ifergan et al., 2024; Wei et al., 2024). Other works used direct observation of activations in order to understand the extent to which transfer occurs (Chen et al., 2024; Zhao et al., 2024). Beyond the usage of fairly complex mechanisms for transfer evaluation that may introduce confounding factors, the main drawback of these white-boxing methods is that they require access to the model’s parameters and therefore can only be applied to open-source models and not to modern proprietary SOTA models.
@@ 5.2 Types of Cross-Lingual Transfer
The first distinction in the literature concerns what is being transferred, specifically the difference between cross-lingual skill transfer and knowledge transfer Cross-lingual skill transfer is the ability to generalize a given skill to unseen languages regardless of the language that was used to learn it. E.g. perfecting multilingual summarization when learning to summarize from English data, or excelling at multilingual instruction following while only exposed to a few languages (Hu et al., 2020; Turc et al., 2021; Malkin et al., 2022; Huang et al., 2023; Shaham et al., 2024). On the other hand, cross-lingual knowledge transfer is the ability to retrieve factual knowledge from one language’s data when queried with any language (Asai et al., 2020; Limkonchotiwat et al., 2022; Mittal et al., 2023; Chua et al., 2024; Litschko et al., 2024).
@@
In addition, methods for evaluation of cross-lingual transfer can also be orthogonally categorized into two broad approaches: fine-tuning on sources, then testing on targets (e.g., Shaham et al., 2024; Limkonchotiwat et al., 2022) vs. evaluation only with the targets (e.g., Malkin et al., 2022; Chua et al., 2024). While the earlier enables higher control over the data and experimental setting, the latter is less expensive and reflects how models behave in the wild.
@@
Lastly, when dealing with cross-lingual knowledge transfer, the knowledge can be either parametric (e.g., Rajaee & Monz, 2024) or contextual (e.g., Chua et al., 2024; Mondshine et al., 2025), a distinction explored by Neeman et al. (2022) in a monolingual setting.
@@
Within this taxonomy of cross-lingual transfer related works, ECLeKTic clearly belongs to methods evaluating parametric knowledge transfer without fine-tuning. In addition, the experiment in subsection 4.4 gradually transition from parametric to contextual knowledge transfer, with the cross-lingual open-book experiment occupying the other end of that spectrum.
@@ 6 Conclusions
In this paper we presented ECLeKTic, a closed-book QA evaluation set for measuring models’ ability to transfer knowledge in their parametric memory across languages. This black-box evaluation is made possible by carefully phrasing questions that target topics highly visible in one language and not in any of the others. Our benchmark allows simple and reliable transfer evaluation which is, for the first time, easily applicable also to API-fenced models. Our results show that cross-lingual knowledge transfer is a difficult task that is far from solved, and that ECLeKTic can be employed to indicate on the progress made towards consistent and inclusive multilingual language models.
@@ 7 Limitations
Although the data ECLeKTic covers varied languages, its coverage is obviously partial. It covers only 6 out of the 10 most spoken languages according to Ethnologue,161616https://www.ethnologue.com/statistics/ and 8 out of the 10 largest Wikipedias in terms of active users.171717https://en.wikipedia.org/wiki/List_of_Wikipedias However, we limited the number of languages to 12 for two reasons: (1) the annotation efforts took about 400 human hours already with 12 languages, and (2) we wanted to keep the benchmark light-weight and because of the exhaustive translation the final number of prompts for evaluation is quadratic in the number of languages.
@@
The limited number of languages also make the conclusions of subsection 4.2 less unequivocal, but the fact the previous works pointed to the same conclusions (Malkin et al., 2022; Mittal et al., 2023; Ifergan et al., 2024) provides some reassurance.
Appendix A Annotation Guidelines
During the creation of ECLeKTic human annotators worked in two stages based on outputs of Gemini (see Appendix B). Here we provide the annotation guidelines there were given in each stage.
A.1 In-Language Verification and Filtering
Purpose
We would like to examine language models’ (LMs) ability to answer questions in a language while seeing little relevant information in that language. For that we scraped a set of Wikipedia pages that appear only in one language and not in the others. The desired set should contain questions like:
Q: “Em qual ano foi criada a Fundação Zoobotânica de Belo Horizonte?” (“In what year was the Zoobotanical Foundation of Belo Horizonte created?”)
We assume that the model has seen the answer to this question in pt-BR, but might not have seen it in other languages, for example ja-JP. We’d like to test the model’s ability to answer this question correctly, when prompted in all other 12 languages. For that it is critical to have a solid set of questions, answers, contexts that are suitable for this purpose.
We would like to ask for your help with creating that dataset.
The Process
The data includes contexts from Wikipedia articles (specifically, the 10 first sentences of each article), and a pair of suggested question and answer based on each context. You need to filter irrelevant examples and modify or rewrite questions and answers that are faulty due to the automatic generation or other reasons.
In general, make sure that the question and its answer are clear, written in proper language. Check whether the question is answerable based on the given article and is faithful to the information conveyed in the article, but also keep in mind that the question should be understandable and answerable even without the given context (for example, by searching the web).
Whenever unsure of how to solve problems with a specific example, feel free to discard the example (i.e., select "to remove" in column D). We want to end up with a high quality dataset, even if slightly smaller.
Specifically, we need you to attend the following potential issues:
Make sure that the Wikipedia article that the example is based on is a contentful article, that is not a disambiguation page or a page dedicated to a list of any kind. If it is, mark the example as “to remove” in column D.
Make sure that the topic of the article and/or the question relates to the speakers of that relevant language. In other words, mark as “to remove” also articles that relate to general topics like science or sociology. Note that a question in, e.g., German about a German biologist counts as a question that relates to the language and not a general scientific question, but a question about natural phenomena – should be removed.
Also note that questions in one language, that refer to topics relevant to another language, are also to be removed. For example, a question in Chinese that refer to a K-pop band that is famous mostly in Korea. -
Make sure that the question is answerable without reference to the content of the article and that it does not include an explicit refernce to the context (e.g., "based on the given paragraph, who was…?"). If this issue is easily fixable, suggest a better phrasing of the question in column E, else – mark it to remove in column D.
Remember that he question are to be translated to other languages so they should be understandable universally. That is to say, that questions that refer to locally-known figures add the relevant information explicitly to the question. For example, when a question is asking about a TV series in India, prepend "the Indian TV series" to its name. You can use Google search to understand whether an entity is locally- or globally-known.
The same kind of disambiguation is needed when referring to certain institutions and other named entities. For example, in a Hebrew question, disambiguate "the supreme court" to "the Israeli supreme court".
A.2 Translation Verification and Modification
Purpose
We would like to examine language models’ (LMs) ability to answer questions in a language while seeing little relevant information in that language. For that we constructed a set of locale-relevant questions, assuming that the model saw information mostly in the language relevant to the locale and not in other languages.
Example: Q: “Em qual ano foi criada a Fundação Zoobotânica de Belo Horizonte?” (“In what year was the Zoobotanical Foundation of Belo Horizonte created?”)
We assume that the model has seen the answer to this question in pt-BR, but might not have seen it in other languages, for example ja-JP. We’d like to test the model’s ability to answer this question correctly, when prompted in all other 12 languages. For that purpose, we automatically translated them to other (target) languages.
We would like to ask for your help with verifying the translations of the questions and with making sure the questions are indeed on subjects related to the language they are presented in.
The Process
The data includes factual questions in 12 languages: German, English, Spanish, French, Hebrew, Hindi, Indonesian, Italian, Japanese, Korean, Portuguese, Chinese (Mandarin), and a translation of each question and answer pair into all other 11 languages, which we’d like to verify, and sometimes slightly fix, if it’s not very natural or doesn’t make sense (see below).
In order to bypass translation between uncommon language pairs, the translation verification will be done in 2 stages:
Verify and correct translations from each of the languages to English, which can be done by a native speaker of each language;
Verify the translation to all other languages, using the now-gold English translation as a reference. Doing so will require every verifier to be knowledgeable only in one language and in English.
Phase 1 Task
In this 1st phase, you only need to verify translations from a given language to English. In cases where the question is generally fine but has some minor issues, you could modify the English translation and, if needed, also the original question. If the fix is impossible or includes a complete rewrite of the question or answer, you can also recommend dropping that question altogether.
Possible Corrections Needed
The automatic translation may have caused multiple problems. Common issues might include unnatural output and missing or added information. In addition, there may be several problems specific to the nature of questions we have at hand, like transliteration of uncommon named entities.
1.
Missing information: In some cases the question is perfectly understandable in the original language, but the translation to English could use a clarification. In these cases modify the original question as well as its translation.
For example, in Spanish, the question: ¿Cuál es la primera de las Siete Palabras? (“what is the first of the Seven Words?”) May be perfectly understandable but in other languages, specifically those with no Christian tradition, it isn’t clear what words are referred to. In this case modify the Spanish original and English translation to clarify that these are the last seven words of Jesus on the cross.
Another example, in Hebrew, the question: my mglm ’t dmwtw /sl db–b‘r h.s‘yr bsdrh ”ml’K m/s.hyt”?! (“Who plays the character of young Dov-Ber in the series ‘Angel of Destruction’?”) is missing an explicit description of the series as an Israeli series, which may make the question somewhat less understandable in other languages. -
2.
Improper transliteration: Sometimes, transliteration depends on some external factors that are missed by the automatic translator. In these cases, the English translation should be amended.
For example, the Chinese question 李显龙的生日是哪一天?refers to 李显龙 who is a Singaporean politician. If the English translation is “when was Li Xianlong born?” it should be amended to reflect the fact that proper transliteration of Chinese names from Singapore does not use the Pinyin system. So the English translation should be modified to “when was Lee Hsien Loong born?” -
3.
Questions that are not locale-specific: If you encounter a question that is not specifically relevant to the source language, please select “Recommend removing” in column D. Such questions can be too general, not specific to any language, for example “How many cm are in an inch?” a question about a Korean actor when the source language is French. If it is unclear if a question is specifically relevant to the source language or not, please do not recommend removing it.
4.
Translatable titles: Names of movies, TV series, books and the like, are a major possible pitfall since they tend to be a translatable phrase and the treatment of uncommon titles is not trivial if they don’t have an official translation. In these cases, we would like to make sure that such official translation indeed does not exist and if so, translate the name as best as possible and make sure it is clearly marked with “…” or «…» or any other way common in the target language.
Phase 2
In the 2nd phase we would like you to verify the translation of the questions from English to your language of expertise. Most questions were translated to English from another language (see phase 1), in these cases you could also take into account the original question, in case you know the source language. In this phase you could only amend or verify the non-English translation. Do not amend the English nor the source language. For common issues that may require amendment of the translation, see phase 1. In cases of severe amendments needed in the source language that were missed in previous iterations, you could also flag the question for removal by selecting “Recommend removing” in column D.
Appendix B Prompts
B.1 ECLeKTic Creation
@@
During the creation of ECLeKTic, LLMs were used for question and answer generation and for the translation of the examples from their respective source languages to all other target languages. Human annotators later verified and fixed their outputs (see Appendix A). We provide prompts for QA generation, translation, LLM as a judge and our ablation prompts in Figure 7, Figure 8, Figure 9 Figure 10 respectively.
Task: Formulate a question in TARGET_LANGUAGE that requires a deep understanding of a given SOURCE_LANGUAGE Wikipedia paragraph. Requirements: * Context-Specific: The question must be answerable solely through information presented within the paragraph, excluding general knowledge or common sense. * Self-Contained: The question should be completely self-explanatory, providing all necessary context within its phrasing. Assume the reader has no access to the paragraph when answering the question. * Single Concrete Factual Detail: - The question should not require multiple answers or involve listing multiple details. - Avoid asking about opinions, interpretations. - In you can’t answer the question, prefer to generate another question. - Focus on extracting a specific, concrete, factual detail that the paragraph directly states. - Be specific: - If you are asking about an entity be clear about it – Use full names for example. - Mention expected granularity: If you are asking about a date, instead of asking "when", ask for a decade, year, month, date etc. If you are asking about a location, instead of asking "where", ask for a country, state, city, street, landmark etc. - Avoid asking questions that their answers are acronyms. Even for non-English examples keep the convention of using the English words "Paragraph", "Response", "question" and "answer" for specifying the parts being generated. Generate only the question and answer. No need to continue with additional examples. Examples: Paragraph: The Great Barrier Reef is the world’s largest coral reef system, composed of over 2,900 individual reefs and 900 islands stretching for over 2,300 kilometers (1,400 mi) over an area of approximately 344,400 square kilometers (133,000 sq mi). The reef is located in the Coral Sea, off the coast of Queensland, Australia. The Great Barrier Reef can be seen from outer space and is the world’s biggest single structure made by living organisms. Response: Question: Where is the Great Barrier Reef located? Answer: Coral Sea, off the coast of Queensland, Australia Paragraph: Die Cazoo Snookerweltmeisterschaft 2023 wurde vom 15. April bis 1. Mai im Crucible Theatre in Sheffield ausgetragen. Mit ihr endete die Saison 2022/23 der World Snooker Tour.[1] Titelverteidiger Ronnie O’Sullivan scheiterte im Viertelfinale gegen Luca Brecel. Der Belgier erreichte das Finale und schlug dort den vierfachen Weltmeister Mark Selby mit 18:15. Brecel ist damit der erste Kontinentaleuropäer, der Weltmeister wurde. In diesem Jahr wurden noch weitere Bestmarken in Bezug auf die 47-jährige „Crucible-Ära“ aufgestellt. Unter anderem übertraf Ronnie O’Sullivan mit seiner 31. Endrundenteilnahme die 30 Teilnahmen von Steve Davis.[2] O’Sullivan erzielte auch sein 200. WM-Century-Break. Zweimal wurde ein Maximum Break erzielt, was es 2008 bereits einmal gegeben hatte; das „perfekte Break“ in einem WM-Finale gelang 2023 erstmals Mark Selby. Response: Question: Gegen wen verlor Ronnie O’Sullivan im Viertelfinale der Snooker-Weltmeisterschaft 2023? Answer: Luca Brecel Paragraph: TEXT Response:
@@
Translate the provided text from IN_LANG to OUT_LANG while maintaining the original meaning and intent. Ensure accuracy and preserve the entities and concepts expressed in the source text. Input in IN_LANG: TEXT Response in OUT_LANG:
@@
Task: Determine if an answer to the question is supported by a given text. Input (in TARGET_LANG): - Text - Question - Answer Single Word Output (in English): - YES: Answer is derived from the text. - NO: Answer is not derived from the text. Text: TEXT Question: QUESTION Answer: PREDICTED_ANSWER Output:
@@
Closed-book QUESTION General hint Answer the following question based on your knowledge in another language. QUESTION Source language name Answer the following question based on your knowledge in {in_lang}. QUESTION Source language title Answer the following question based on your knowledge in {in_lang} about {original_title}. QUESTION Cross-lingual open-book Context: SOURCE_LANG_CONTEXT QUESTION
Appendix C Per-Language Breakdown - Overall Success